When working with exponents, is there any difference in operations if the exponents are whole numbers or fractions? In what sense do exponents and radicals behave exactly the same way? Provide examples
Answer by stanbon(75887)

When working with exponents, is there any difference in operations if the exponents are whole numbers or fractions?
There is no difference.
==================================
In what sense do exponents and radicals behave exactly the same way?
[sqrt(a)*sqrt(b) = sqrt(ab)
----
a^(1/2) * b^(1/2) = (ab)^(1/2)
-----------------
Cheers,
Stan H.